The function f defined by f(x) = tan(2x) is a periodic function with period π. ii) The function 𝑓:𝑹 → 𝑹, defined by 𝑓(𝑥) = 1 − |𝑥| is differentiable at x=1. iii) The function 𝑓: [3,4] → 𝑹 defined by 𝑓(𝑥) = 𝑥 ଶ − 𝑥 is monotonic in its domain. iv) Every continuous function is differentiable. v) Every curve over R has a point of inflection. 


1
Expert's answer
2022-01-24T16:07:52-0500

(i)

Function f(x) = tanx is a periodic function with period "\pi".

So, period of tan2x is "\pi"/2. This is because period of tanax is "\pi"/a.


The given statement is false.


(ii)

Lets graph f(x) = 1 - |x|




Clearly, at x = 1, there is only tangent. So, the function is differentiable at x = 1.


The given statement is true.


(iii)

"f(x)=x^2-x"




A monotonic function is a function which is either entirely nonincreasing or nondecreasing.

The given function is nondecreasing in [3,4], so it is monotonic in its domain.


So the given statement is correct.


iv)

for example:

"f(x)=|x|" is continuous but not differentiable at x = 0.


statement is false


v)

if the second derivative can equal zero, the original function has a point of inflection

for example:

"y=e^x"

"y''=e^x" cannot equal zero, so the curve has not a point of inflection


statement is false